OpenStudy (anonymous):

Solve the equation of (¼)log6 (a – 3) – log6 3 = 0.

OpenStudy (anonymous):

¼)log6(a – 3) – log6(3) = 0 log6[(a – 3)^¼] = log6(3) So (a – 3)^¼ = 3 a - 3 = 3^4 = 81 a = 81 + 3 = 84